# Initialization using Update Approximation is a Silver Bullet for Extremely Efficient Low-Rank Fine-Tuning

### v1

@@ 3.4 Initialization using update approximation
Let us first consider the objective in fine-tuning, the primary goal is to update weights to better suit the target task and domain. The initial gradient steps are particularly informative, as they indicate the direction of desired adaptation. We propose leveraging this insight by using the first update step from full fine-tuning to initialize our low-rank matrices.
This approach offers two key advantages: First, it ensures the low-rank space captures the most relevant subspace for the target task rather than relying on pre-trained weight properties. Second, since $A$ and $B$ matrices will remain fixed in LoRA-XS, initializing them to span the subspace of early adaptation increases the likelihood of capturing useful updates throughout training. This would also ensure that the final update is learnt in the correct subspace, of which we have no apriori information besides the first step of full fine-tuning.
In essence, our method can be summarized as follows: Set an initialization which best approximates the first step of full fine-tuning. Formally, given a full fine-tuning update $\Delta W$, our initialization should satisfy:
@@
The first step of full fine-tuning, for Adam-based optimizers such as AdamW, for a sample $x_{i}$ is:
@@
Using a single sample may lead to noisy estimates. Instead, we compute a more stable initialization by averaging gradients over a subset of the training data:
$\Delta W_{avg}=-\eta\textbf{sign}(\sum_{i=0}^{n\leq|\mathbb{X}|}\nabla_{W}\mathcal{L}(W_{0},x_{i}))\dots x_{i}\in\mathbb{X}$ (7)
This averaged gradient better captures the general direction of adaptation required for the target task while being less sensitive to individual sample variations. We can then use truncated SVD to obtain a low-rank approximation of $\Delta W_{\text{avg}}$, and express it as $sBRA$. There exist infinite combinations of $B$ and $A$ which can obey the above relationship. For instance, we can initialize $B$ and $A$ as $US$ and $V^{T}$ and keep $R$ as $I/s$. This initialization is equivalent to the $B$ and $A$ initialization in LoRA-XS but by approximating the update rather than the pre-trained matrix. We note that the above process can be computed for any optimizer, by approximating the corresponding first step. We compute this specifically for AdamW since we use it.
@@ 3.5 Hyperparameter independence
The hyperparameter $\alpha$ is used in LoRA and subsequent decomposition based models to tackle the issue of instability caused to improper scaling of the updates. The gradient scaling is accounted for, by adding a hyperparameter to normalize the updates. The importance of scaling is well documented in methods like rank stabilization (Kalajdzievski, 2023). However, the full fine-tuning gradient $g$ needs no such tuning. We claim that approximating the full fine-tuning gradient removes the need for introducing a scaling factor.
@@
Appendix A.5
The hyperparameter independence of the equivalent gradient eliminates the need for manual gradient scaling. Updates to weight matrix $W$ depend solely on this gradient (modulo learning rate), making any additional scaling redundant.
This can be understood by examining the relationship with full fine-tuning gradient $g$. Since $g$ is naturally scaled for optimal weight updates, and our method approximates $g$ in a constrained subspace, the equivalent gradient inherits appropriate scaling automatically. Notably, this property is unique to our gradient approximation approach and does not hold for standard LoRA-XS training.
@@ 3.6 LoRA-SB: Update approximation initialization is a silver bullet
The solutions discussed above independently address the gradient approximation and initialization problems, while also providing hyperparameter independence. Our proposed method, LoRA-SB, elegantly combines these solutions through a simple initialization strategy.
Our initialization strategy is derived from approximating the first step of full fine-tuning:
$U,S,V^{T}\leftarrow\textbf{SVD}(\Delta W_{avg})$ (8)
@@
By the Eckart-Young theorem (Eckart & Young, 1936; Mirsky, 1960), this gives the optimal rank-$r$ approximation of the initial full fine-tuning update. where $U$, $S$, $V$ are obtained from truncated SVD of the averaged first update $\Delta W_{\text{avg}}$.
This initialization leads to several key advantages that address the problems identified earlier.
Simplified Gradient Optimization. This initialization ensures $B_{\text{init}}$ and $A_{\text{init}}$ form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$ respectively, leading to $B^{T}B=AA^{T}=I$. With fixed $B$ and $A$ matrices being orthonormal, the optimal update step derived in Equation 3 simplifies to:
$g^{R}=\dfrac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{LoRA-XS}(AA^{T})^{-1}=\frac{1}{s^{2}}g^{R}_{LoRA-XS}$ (12)
This eliminates the need for complex matrix inversions during training.
Optimum Update Approximation. Our initialization guarantees that the first update optimally approximates the full fine-tuning weight updates:
$sB_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$ (13)
By the Eckart-Young theorem, this gives the optimal rank-$r$ approximation of the initial full fine-tuning update.
Hyperparameter Independence. As shown in Theorem 5, when gradient approximation is applied with orthonormal $B$ and $A$, the hyperparameter $s$ can be set to 1, resulting in the following:
@@
This demonstrates that our initialization guarantees optimal gradient approximation at every step, without requiring any scaling factor!
@@
Thankfully, we don’t have to choose between the two. For SGD, we prove that setting $B_{init}$ and $A_{init}$ using Equations 8-11, results in the first update of LoRA-XS to best approximate the direction of the update of full fine-tuning.
@@
If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
@@
Appendix A.6
LoRA-SB Advantages over LoRA. Notably, many of the properties described above are not achievable with standard LoRA methods. Even if $B$ and $A$ are initialized as orthonormal in LoRA, subsequent updates do not preserve this property because $B$ and $A$ are trainable. This results in several challenges:
  • Potential instability of $(B^{T}B)^{-1}$ and $(AA^{T})^{-1}$, as they are not guaranteed to remain non-singular during training.
@@
  • Repeated re-computation of $B^{T}B$ and $AA^{T}$ is required at each optimizer step for accurate gradient approximation.
LoRA-SB’s simple initialization strategy solves multiple problems simultaneously, offering a robust and efficient approach to low-rank adaptation.
Algorithm. We present a PyTorch-like implementation of our method in Algorithm 1. To optimize GPU memory usage during initialization, we hook into the backward pass of PyTorch and compute the gradient for one layer at a time, immediately discarding the computed gradients (Lv et al., 2024; Wang et al., 2024a). This ensures that memory usage remains at $O(1)$, independent of the number of layers. This approach keeps memory consumption under control, ensuring it never exceeds the requirements of subsequent LoRA-SB fine-tuning. For very large batch sizes, one can also use gradient accumulation and quantization to further manage memory usage within the specified limits.
@@
$\text{{model}}\leftarrow$ AutoModel(base_model)
@@
$\text{{trainer}}\leftarrow$ Trainer(sb_model, ...)
$\text{{trainer}}.$train()
@@ 4 Experiments
We implement all models using PyTorch (Paszke et al., 2019) and the popular HuggingFace Transformers library (Wolf et al., 2020). We run all experiments on a single NVIDIA A6000 GPU and report results as the average of three random seeds. To save memory, we initialize base models in torch.bfloat16 precision. Appendix D provides detailed information on the datasets used.
We compute the update approximation using only $\mathbf{1/1000}$ of each dataset’s total number of samples. This ensures that the additional training time overhead is minimal and has a negligible effect on overall efficiency. We provide a benchmark of this overhead in Section 5. These samples are randomly selected from the dataset’s training set in each run.
@@ 4.1 Arithmetic Reasoning
Details. We fine-tune Mistral-7B (Jiang et al., 2023) and Gemma-2 9B (Team et al., 2024) on 50K samples from the MetaMathQA (Yu et al., 2024) dataset and evaluate them on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks. Evaluation focuses solely on the final numeric answer. We apply LoRA modules to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$. Detailed hyperparameter settings are provided in Appendix C.
Results. We present the results for both models in Table 1. LoRA-SB significantly outperforms LoRA-XS across all parameter settings. Notably, LoRA-SB surpasses the performance of LoRA ($r=32$) while using 40x fewer trainable parameters for Mistral-7B and 90x fewer for Gemma-2 9B at ranks $r=96$ and $r=64$, respectively.
We present training loss curves comparing LoRA-SB and LoRA-XS in Figure 2. Thanks to superior initialization, LoRA-SB starts with a lower initial loss compared to LoRA-XS. Additionally, due to optimal gradient approximation, LoRA-SB maintains a consistently better loss curve throughout training and converges to a superior final value.
Table 1: Accuracy comparison of fine-tuning methods on Mistral-7B and Gemma-2 9B across the arithmetic reasoning benchmarks GSM8K and MATH, after training on MetaMathQA. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold.
@@ 4.2 Commonsense Reasoning
Details. We fine-tune Llama-3.2 3B (Dubey et al., 2024) on CommonSense170K, a dataset that aggregates eight commonsense reasoning tasks (Hu et al., 2023). We evaluate the model’s performance on each dataset individually, which include BoolQ (Clark et al., 2019), SIQA (Sap et al., 2019), PIQA (Bisk et al., 2020), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2021), and HellaSwag (Zellers et al., 2019). Each example is framed as a multiple-choice question where the model generates the correct answer without explanations. We use the prompt template from (Hu et al., 2023). LoRA modules are applied to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$. Hyperparameter details are provided in Appendix C.
Results. We present the results in Table 2. LoRA-SB consistently outperforms LoRA-XS across all parameter settings. Additionally, LoRA-SB ($r=96$) exceeds the performance of LoRA ($r=32$) with 27x fewer trainable parameters.
Table 2: Accuracy comparison of fine-tuning methods on Llama-3.2 3B across eight commonsense reasoning datasets. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold.
@@
LoRA $32$ $48.63$ M $70.03$ $\mathbf{85.20}$ $79.12$ $90.71$ $82.24$ $86.91$ $74.32$ $\mathbf{81.87}$ $81.44$
@@
LoRA-SB $64$ $0.80$ M $68.35$ $84.55$ $79.94$ $\mathbf{91.68}$ $83.03$ $87.84$ $74.83$ $80.12$ $81.29$
LoRA-SB $96$ $1.81$ M $\mathbf{70.34}$ $84.76$ $\mathbf{80.19}$ $91.62$ $\mathbf{84.61}$ $\mathbf{87.92}$ $\mathbf{74.74}$ $81.20$ $\mathbf{81.92}$
@@ 4.3 Natural Language Understanding
Details. We fine-tune RoBERTa-large (Liu et al., 2019) on GLUE, a sequence classification benchmark that contains several datasets, covering domains like sentiment analysis and natural language inference. The datasets we evaluate on are: CoLA (Warstadt et al., 2019), RTE, MRPC (Dolan & Brockett, 2005), SST-2 (Socher et al., 2013), QNLI (Rajpurkar et al., 2018), and STS-B (Cer et al., 2017). LoRA modules are applied only to the self-attention layers, following the configuration in the original LoRA paper (Hu et al., 2021), with ranks $r=\{8,16,24\}$. Detailed experimental settings are available in Appendix C.
Results. The results are shown in Table 3. LoRA-SB consistently outperforms LoRA-XS across all parameter configurations. Notably, LoRA-SB ($r=24$) outperforms LoRA ($r=8$) with 39x lesser trainable parameters.
Table 3: Comparison of fine-tuning methods on RoBERTa-large across the GLUE benchmark datasets. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold. We use Pearson correlation for STS-B, Matthew’s correlation for CoLA, and accuracy for others. Higher is better for each metric.
@@
LoRA 88 $2162.69$ K $68.02$ $82.98$ $90.05$ $91.43$ $93.42$ $95.98$ $86.98$
LoRA-XS 88 $6.14$ K $61.07$ $75.23$ $86.21$ $89.29$ $92.44$ $94.72$ $83.16$
LoRA-XS $16$ $24.57$ K $63.32$ $79.06$ $86.28$ $90.36$ $93.89$ $95.76$ $84.70$
@@
LoRA-SB 88 $6.14$ K $63.57$ $78.43$ $88.72$ $90.59$ $92.95$ $95.07$ $84.88$
LoRA-SB $16$ $24.57$ K $64.36$ $82.31$ $89.71$ $91.24$ $93.89$ $95.87$ $86.23$
LoRA-SB $24$ $55.20$ K $\mathbf{68.28}$ $\mathbf{83.03}$ $\mathbf{90.12}$ $\mathbf{91.65}$ $\mathbf{93.75}$ $\mathbf{96.11}$ $\mathbf{87.16}$

### v3

@@ 3.4 Initialization using update approximation
In FT, the primary goal is to update weights to better suit the target task. The initial gradient steps are particularly informative, as they indicate the direction of desired adaptation. We leverage this insight by using the first update step from full FT for initialization.
This approach offers two key advantages. First, it ensures the low-rank space captures the most relevant subspace for the target task rather than relying on pre-trained properties. Second, since $A$ and $B$ are fixed, initializing them to span the subspace of early adaptation increases the likelihood of capturing useful updates throughout training. This also ensures that the final update is learnt in the correct subspace, of which we have no apriori information besides the first full FT step. Our method is summarized as: set such initialization that best approximates the first step of full FT. Given a full FT update $\Delta W_{first-step}$, our initialization satisfies:
@@
The first step of full FT, for Adam-based optimizers such as AdamW, for sample $x_{i}$ is:
@@
However, the usage of a single sample may lead to noisy estimates. Instead, we compute a more stable initialization by averaging gradients over a subset of the training data:
$\Delta W_{avg}=-\eta\textbf{sign}(\sum_{i=0}^{n\leq|\mathbb{X}|}\nabla_{W}\mathcal{L}(W_{0},x_{i})),\quad x_{i}\in\mathbb{X}$ (7)
This better captures the general direction of adaptation required for the target task while being less sensitive to individual sample variations. We can then use truncated SVD to obtain a low-rank approximation of $\Delta W_{\text{avg}}$, and express it as $sBRA$. There exist infinite combinations of $B$ and $A$ which can obey this relationship. For instance, we can initialize $B$ and $A$ as $US$ and $V^{\top}$ and keep $R$ as $I/s$. This is equivalent to the $B$ and $A$ initialization in LoRA-XS but by approximating the update rather than the pre-trained matrix. We note that the above process can be computed for any optimizer, by approximating the corresponding first step. We compute this specifically for AdamW since we use it.
@@ 3.5 Hyperparameter independence
The hyperparameter $\alpha$ is used in LoRA and other decomposition-based method to tackle the issue of instability caused to improper scaling of the updates. The gradient scaling is accounted for, by adding a hyperparameter to normalize the updates. The importance of scaling is shown in methods like rank stabilization (Kalajdzievski, 2023). However, the full FT gradient $g$ needs no such tuning. We claim that approximating the full FT gradient removes the need for introducing a scaling factor, as shown in Theorem 5.
@@
See Appendix A.5. ∎
The hyperparameter independence of the equivalent gradient eliminates the need for manual gradient scaling. Updates to $W$ depend solely on this gradient (modulo learning rate), making any additional scaling redundant. This can be understood by examining the relationship with the full FT gradient $g$. Since $g$ is naturally scaled for optimal weight updates, and our method approximates $g$ in a constrained subspace, the equivalent gradient inherits appropriate scaling automatically. This property is unique to our gradient approximation approach and does not hold for standard LoRA-XS.
@@ 3.6 LoRA-SB: Update approximation initialization is a silver bullet
The solutions discussed above independently address the gradient approximation and initialization problems, while also providing hyperparameter independence. Our proposed method, LoRA-SB, elegantly combines these solutions through a simple initialization strategy, derived from approximating the first step of full FT:
$U,S,V^{\top}\leftarrow\textbf{SVD}(\Delta W_{avg})$ (8)
@@
By the Eckart-Young theorem (Eckart & Young, 1936; Mirsky, 1960), this gives the optimal rank-$r$ approximation of the full FT update. where $U$, $S$, $V$ are obtained from truncated SVD of the averaged first update $\Delta W_{\text{avg}}$. This initialization leads to several key advantages.
Simplified Gradient Optimization. Our initialization ensures $B_{\text{init}}$ and $A_{\text{init}}$ form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$ respectively, leading to $B^{\top}B=AA^{\top}=I$. With fixed $B$ and $A$ matrices being orthonormal, the need for complex matrix inversions during training is eliminated, , as the optimal update step, derived in Equation 3, simplifies to:
$g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}=\frac{1}{s^{2}}g^{R}_{LoRA-XS}.$
Optimal Update Approximation. Our initialization guarantees that the first update optimally approximates the full FT weight updates:
$sB_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}.$
By the Eckart-Young theorem, this gives the optimal rank-$r$ approximation of the initial full FT update.
Hyperparameter Independence. As shown in Theorem 5, when gradient approximation is applied with orthonormal $B$ and $A$, the hyperparameter $s$ can be set to 1, resulting in:
@@
This demonstrates that our initialization guarantees optimal gradient approximation at every step, without requiring any scaling factor.
@@
Thankfully, we don’t have to choose between the two. For SGD, we prove that setting $B_{init}$ and $A_{init}$ using Equations 8-11, results in the first update of LoRA-XS to best approximate the direction of the update of full FT (Theorem 6).
@@
If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
@@
See Appendix A.6. ∎
LoRA-SB Advantages over LoRA. Many properties described above are not achievable with standard LoRA methods. Even if $B$ and $A$ are initialized as orthonormal in LoRA, subsequent updates do not preserve this property because $B$ and $A$ are trainable. This results in several challenges:
  • Potential instability of $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$, as they are not guaranteed to remain non-singular during training.
@@
  • Repeated re-computation of $B^{\top}B$ and $AA^{\top}$ is required at each optimizer step for accurate gradient approximation.
Algorithm. We present a PyTorch-like implementation of our method in Algorithm 1. To optimize GPU memory usage during initialization, we hook into the backward pass of PyTorch and compute the gradients layerwise, immediately discarding the computed gradients (Lv et al., 2024; Wang et al., 2024a). This ensures $O(1)$ memory usage, independent of the number of layers, keeping memory consumption well within manageable limits and ensuring it does not exceed the memory requirements of subsequent LoRA-SB FT. For large batch sizes, memory usage can be further optimized through gradient accumulation and quantization.
@@
[…]
@@
[…]
@@ 4 Experiments
We use PyTorch (Paszke et al., 2019) and the HuggingFace Transformers library (Wolf et al., 2020) for our implementations. We run all experiments on a single NVIDIA A6000 GPU and report results as the average of three random seeds. To save memory, we initialize base models in torch.bfloat16 precision. Appendix D provides detailed information on the datasets used. We compute the update approximation using only $\mathbf{1/1000}$ of each dataset’s total number of samples. This ensures that the additional training time overhead is minimal and has a negligible effect on overall efficiency. We provide a benchmark of this overhead in Section 5. The samples are randomly selected from the training set in each run. Detailed hyperparameter settings for all tasks are given in Appendix C.
Baselines. We compare LoRA-SB against full FT, LoRA, LoRA-XS, rsLoRA, and PiSSA. rsLoRA introduces a rank-scaled stabilization factor ($\alpha/\sqrt{r}$) to enhance stability, while PiSSA updates only the principal components of the pre-trained weight $W$ and freezes the residuals.
@@ 4.1 Arithmetic Reasoning
Table 1: Accuracy comparison of FT methods on Mistral-7B and Gemma-2 9B across the arithmetic reasoning benchmarks GSM8K and MATH, after training on MetaMathQA. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold.
@@
Details. We fine-tune Mistral-7B (Jiang et al., 2023) and Gemma-2 9B (Team et al., 2024) on 50K samples from the MetaMathQA (Yu et al., 2024) dataset and evaluate them on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks. We apply LoRA modules to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
Results. We present results in Table 1. LoRA-SB significantly outperforms LoRA-XS across all settings. Notably, LoRA-SB outperforms LoRA-based methods ($r=32$) while using 40x fewer trainable parameters for Mistral-7B and 90x fewer for Gemma-2 9B at ranks $r=96$ and $r=64$, respectively. We present training loss curves comparing LoRA-SB and LoRA-XS in Figure 2. Thanks to superior initialization, LoRA-SB starts with a lower initial loss compared to LoRA-XS. Additionally, due to optimal gradient approximation, LoRA-SB maintains a consistently better loss curve throughout and converges to a superior final value.
@@ 4.2 Commonsense Reasoning
Details. We fine-tune Llama-3.2 3B (Dubey et al., 2024) on CommonSense170K, a dataset with eight commonsense reasoning tasks (Hu et al., 2023). We evaluate the model’s performance on each dataset individually, which include BoolQ (Clark et al., 2019), SIQA (Sap et al., 2019), PIQA (Bisk et al., 2020), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2021), and HellaSwag (Zellers et al., 2019). LoRA modules are applied to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
Results. We present the results in Table 2. LoRA-SB consistently outperforms LoRA-XS across all settings. In addition, LoRA-SB ($r=96$) outperforms LoRA-based methods ($r=32$) with 27x fewer trainable parameters.
Table 2: Accuracy comparison of FT methods on Llama-3.2 3B across eight commonsense reasoning datasets. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold.
@@
LoRA $32$ $48.63$ M $70.03$ $85.20$ $79.12$ $90.71$ $82.24$ $86.91$ $74.32$ $\mathbf{81.87}$ $81.30$
rsLoRA $32$ $48.63$ M $69.81$ $85.05$ $78.92$ $90.45$ $82.02$ $86.71$ $74.18$ $81.72$ $81.11$
PiSSA $32$ $48.63$ M $70.12$ $\mathbf{85.42}$ $79.44$ $90.88$ $82.68$ $87.23$ $74.61$ $81.79$ $81.52$
@@
LoRA-SB $64$ $0.80$ M $68.35$ $84.55$ $79.94$ $\mathbf{91.68}$ $83.03$ $87.84$ $\mathbf{74.83}$ $80.12$ $81.29$
LoRA-SB $96$ $1.81$ M $\mathbf{70.34}$ $84.76$ $\mathbf{80.19}$ $91.62$ $\mathbf{84.61}$ $\mathbf{87.92}$ $74.74$ $81.20$ $\mathbf{81.92}$
@@ 4.3 Natural Language Understanding
Details. We fine-tune RoBERTa-large (Liu et al., 2019) on GLUE, a popular language understanding classification benchmark that contains several datasets. The datasets we evaluate on are: CoLA (Warstadt et al., 2019), RTE, MRPC (Dolan & Brockett, 2005), SST-2 (Socher et al., 2013), QNLI (Rajpurkar et al., 2018), and STS-B (Cer et al., 2017). LoRA modules are applied only to the self-attention layers, with ranks $r=\{8,16,24\}$.
Results. The results are shown in Table 3. LoRA-SB consistently outperforms LoRA-XS across all configurations. Additionally, LoRA-SB ($r=24$) outperforms LoRA-based methods ($r=8$) with 39x lesser trainable parameters.
Table 3: Comparison of FT methods on RoBERTa-large across the GLUE benchmark datasets. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold. We use Pearson correlation for STS-B, Matthew’s correlation for CoLA, and accuracy for others. Higher is better for each metric.
@@
LoRA $8$ $2162.69$ K $68.02$ $82.98$ $90.05$ $91.43$ $93.42$ $95.98$ $86.98$
rsLoRA $8$ $2162.69$ K $67.87$ $82.84$ $89.97$ $91.30$ $93.29$ $95.87$ $86.85$
PiSSA $8$ $2162.69$ K $68.22$ $\mathbf{83.14}$ $90.10$ $91.59$ $93.55$ $96.03$ $87.10$
LoRA-XS $8$ $6.14$ K $61.07$ $75.23$ $86.21$ $89.29$ $92.44$ $94.72$ $83.16$
LoRA-XS $16$ $24.57$ K $63.32$ $79.06$ $86.28$ $90.36$ $93.69$ $95.76$ $84.70$
@@
LoRA-SB $8$ $6.14$ K $63.57$ $78.43$ $88.72$ $90.59$ $92.95$ $95.07$ $84.88$
LoRA-SB $16$ $24.57$ K $64.36$ $82.31$ $89.71$ $91.24$ $\mathbf{93.89}$ $95.87$ $86.23$
LoRA-SB $24$ $55.20$ K $\mathbf{68.28}$ $83.03$ $\mathbf{90.12}$ $\mathbf{91.65}$ $93.75$ $\mathbf{96.11}$ $\mathbf{87.16}$